Show and check work.
-3(4-x)=5-(x+1)

Answer by checkley77(12844)

-3(4-x)=5-(x+1)
-12+3x=5-x-1
3x+x=5-1+12
4x=16
x=16/4
x=4 ans.
Proof:
-3(4-4)=5-(4+1)
-3*0=5-5
0=0